OpenStudy (anonymous):

Please help me :( The population of a certain geographic region is approximately 133 million and grows continuously at a relative growth rate of 1.59%. What will the population be in 6 years? Compute the answer to three significant digits.

jimthompson5910 (jim_thompson5910):

You use the formula A = Pe^(rt) where A = unknown (solving for it)(population in t years) P = 133 (population now -- in millions) r = 0.0159 (growth rate) t = 6 (# of years) A = Pe^(rt) A = 133*e^(0.0159*6) A = 133*e^(0.0954) A = 133*1.10009880665 A = 146.31314128445 A = 146 (round to three significant digits) So the population will be roughly 146 million in 6 years